Westpoint Company is considering an investment project that generates a cash flow of $1,350,000 next year if the economy is favorable but generates only $620,000 if the economy is unfavorable. The probability of favorable economy is 60% and of unfavorable economy is 40%. The project will last only one year and be closed after that. The cost of investment is $880,000 and Westpoint Company plans to finance the project with $180,000 of equity and $700,000 of debt. Assuming the discount rates of both equity and debt are 0%. What are the expected cash flows to Westpoint Company s creditors and shareholders if the company invests in the project? $668,000 to creditors and $0 to shareholders $584,000 to creditors and $0 to shareholders $584,000 to creditors and $360,000 to shareholders $668,000 to creditors and $390,000 to shareholders $700,000 to creditors and $520,000 to shareholders
Answer #1

Discount rate of Equity and debt is is 0%. so wacc is also 0%.      
Present value = Expected cash flow in year 1/(1+Discount rate)^year      
Expected cash flow is $1350000, if economy is favourable. So Present value =1350000/(1+0%)^1 = 1350000.

Debt is issued for $700000. First debt will be paid.

so Expected cash flow to debt if economy is favourable will be $700000       
Cash flow remaining to Equity shareholders = Present value of cash flow - Debt payment      
1350000-700000      
=$650000


Expected cash flow is $620000, if economy is unfavourable. So Present value =620000/(1+0%)^1 = 620000.

Payment priority is debt. So whole $620000 will be paid to creditors or Debt holders.

Expected cash flow to debt if economy is unfavourable will be $620000      
Cash flow remaining to Equity shareholders = Present value of cash flow - Debt payment      
620000-620000      
=0      
      
Expected cash flow to Equity next year = (Cash flow in favourable *probability)+(Cash flow in unfavorable * probability)      
(650000*60%)+(0*40%)      
=390000      
So Expected cash flow to Equity is $390000      
      
Expected cash flow to Debt or creditors next year = (Cash flow in favourable *probability)+(Cash flow in unfavorable * probability)      
(700000*60%)+(620000*40%)      
=668000      


So Expected cash flow to creditors is $668000 and to Equity is $390000